Identify the solution(s) of the system of equations, if any.
3x+2y=4 and 5x-4y=3
My options for answers are
(1,.5)
(.5, 1)
(5, 5.5)
(5/11, -2/11)
I asked for help on a question similar to this only the 2nd problem was already in the y=mx+b format. Not sure if this would be done the same way?

Answer by richwmiller(17219)

3x+2y=4
5x-4y=3
double the first equation
6x+4y=8
5x-4y=3
add them
11x=11
x=1
the only option with x=1 is the first
(1,.5)
3+2y=4
2y=1
y=1/2=.5